Justify that


lim 1/(5-x)^2
x tend to5 = ∞
1
Expert's answer
2020-03-09T14:10:11-0400

"\lim\limits _{x\mapsto 5}\frac{1}{(5-x)^2}=\infty\\
x_0=5, f(x)=\frac{1}{(5-x)^2}."

It must be shown that

"\forall k>0\quad \exist\delta(k)>0, \quad |x-x_0|<\delta"

performed

"|f(x)|>k."

Consider

"|x-5|<\delta"

then

"\frac{1}{|x-5|}>\frac{1}{\delta}"

because

"\delta>0, \quad |x-5|>0."

Consider

"|f(x)|=\left|\frac{1}{(5-x)^2}\right|=\frac{1}{|(x-5)^2|}=\\
=\frac{1}{|x-5|^2}>\frac{1}{\delta^2}>k\implies\\
\frac{1}{\delta^2}>k \implies \delta^2<\frac{1}{k} \implies\\
 \delta<\frac{1}{\sqrt{k}}."

So at "\delta<\frac{1}{\sqrt{k}}" have "\left|\frac{1}{(5-x)^2}\right|>k" ,

which means that

"\lim\limits _{x\mapsto 5}\frac{1}{(5-x)^2}=\infty."